OpenStudy (anonymous):

The period of the moon is approximately 27.2 days. Determine the radius of the moon's orbit and the orbital speed of the moon.

OpenStudy (ivancsc1996):

Well which equations do you know about orbital movement?

OpenStudy (anonymous):

\[V =\sqrt{Gm divr}\]

OpenStudy (anonymous):

\[g=Gm/r^2\]

OpenStudy (ivancsc1996):

Ok so the only real problem is finding the radius of the moon's orbit. But I can't find it with any of those equations. Don't they give you the gravitational force between the earth and the moon or a equation involving the centripetal force?

OpenStudy (anonymous):

\[Fc=mv^2/r\] \[Fg=Gmm/r^2\]

OpenStudy (ivancsc1996):

Well \[F _{c}=F _{g}=\frac{ mv ^{2} }{ r }=\frac{ Gm _{1}m _{2} }{ r ^{2} } \rightarrow r=\frac{ Gm _{1} }{ v ^{2} }\]But that is no different from the equation in the previous comments. I steel need something more.